# VimTS: A Unified Video and Image Text Spotter for Enhancing the Cross-domain Generalization

### v3

@@ Abstract
Unified; Synergy; Video Text Spotting; Text Detection; Text Recognition; Text Tracking
@@ 1 Introduction
Figure 1: Fig. (a) and (b) are two types of cross-domain text spotting, including image-to-image and image-to-video. TT represents the TotalText. IC15 represents the ICDAR2015 [8]. IC13 represent the ICDAR2013 video [9]. TT represents the TotalText [10]
1) In image-level cross-domain text spotting, the challenge lies in the varying styles, fonts, and backgrounds across different benchmarks, requiring models to generalize well beyond their training data. When the training set and the test set belong to different benchmarks, the model usually performs poorly, as demonstrated in  [7]. Furthermore, the differences in annotation formats across various datasets bring further challenges. For instance, Total-Text [10] and ICDAR2015 [8] utilize word-level annotations, while CTW1500 [11] employs line-level annotations. Integrating datasets with distinct annotation formats simultaneously proves challenging. This disparity not only complicates the training process but also hampers model performance, as the model struggles to adapt to divergent annotation schemes.
Figure 2: Applying static text spotting methods to videos, even those with minimal motion, results in inadequate performance regarding both bounding box recall and recognition accuracy. $T_{n}$ represents the $n_{th}$ frame of a video.
@@ 2.1 Scene Text Spotting
Scene text spotting has seen a significant shift from conventional methodologies to more modern end-to-end learning techniques. Traditional approaches, as highlighted in research such as [16], often faced challenges related to inference inefficiencies and error accumulation. Pioneering a more unified perspective, Li et al. [17] integrated detection and recognition into a unified framework, creating a comprehensive framework, which predominantly focused on horizontal texts.
@@ 2.2 Video Text Spotting
Compared to static text spotting, there have been few recent advancements in video text spotting due to its challenges. Yin et al. [31] surveyed text detection [32], tracking [33], and recognition in video. Wang et al. [34] developed an end-to-end system for video text spotting, utilizing hand-crafted post-processing for tracking. Subsequently, Cheng et al. [35] integrated a spatial-temporal video text detector with a text recommender to select high-quality text and perform recognition once. Its improved version, FREE [36], adopted an end-to-end trainable framework for faster and robust video text spotting. Nguyen et al. [37] performed multi-scale character detection and exploited temporal redundancy to remove false positives. Rong et al. [38] proposed a framework for scene text recognition across multiple frames using feature representation and a conditional random field model. Wu et al. [39] developed a dual decoder framework for detection and tracking, utilizing object features and IoU match. Inspired by the MOTR [40], Wu et al. [41] unified text detection and tracking into a single decoder, treating video text spotting as a long-range temporal variation problem. To further enhance the performance of the recognizer, VLSpotter [42] introduced a super-resolution module and language model for video text spotting.
@@ 2.3 Domain Adaptation
Domain adaptation, a burgeoning field of study, addresses the challenges posed by domain shift and dataset bias. It has been studied extensively in many fields [43, 44, 45]. Researchers also proposed several cross-domain approaches in OCR. For instance, Chen et al. [46] propose a pixel-level domain adaptation block and an image-level domain adaptation block for text detection to reduce the domain shift by introducing adversarial training. Wu et al. [47] attempt to transfer knowledge from synthetic data to real data for scene text detection. They employ text instance alignment to facilitate learning domain-invariant features adversarially, and utilize text self-training to reduce the impact of inaccurate pseudo-labels, hence minimizing false positives (FPs) and false negatives (FNs). Additionally, Chen et al. [48] develop a self-training framework aimed at automatically extracting the pseudo-labels of challenging text instances from unannotated videos or images. In another work, Yu et al. [49] introduce a cross-modal interaction mechanism, integrating vision-language information from CLIP [50] to improve the cross-domain performance of scene text detectors. The above methods are designed for scene text detection.
For scene text recognition, Zhang et al. [51] introduced SSDAN, a sequence-to-sequence domain adaptation network. SSDAN utilizes a gated attention similarity unit to dynamically align the distribution of source and target sequence data within a character-level feature space. GA-DAN [52] develops a geometry-aware domain adaptation network for both text detection and recognition. It effectively handles cross-domain shifts by simultaneously modeling changes in both geometric and appearance spaces, enabling realistic image conversion between domains with diverse characteristics. In recent years, there have also been ways to start looking at cross-domain aspects of text spotting. Expanding upon their prior work [49], Yu et al. [7] propose FastTCM-CR50 as a new backbone for scene text detection and spotting, to improve the cross-domain capability of existing methods. FastTCM-CR50 incorporates an implicit image condition and bimodal similarity matching module, empowering the CLIP text encoder to conduct offline computations during inference. As a result, it achieves superior performance and reduces inference time.
@@ 3 Methodology
In this paper, we introduce VimTS, a unified framework designed to leverage the synergy between various tasks. Through this unified framework, this approach aims to improve the generalization ability of the text spotter. The overall architecture is illustrated in Fig. 3. Inspired by [14], we employ a set of task-aware queries to represent various tasks. Initially, image features are obtained via a feature extraction process comprising ResNet50, the receptive enhancement module (REM), and the Transformer encoder. Similar to [14], these features are then used to generate task-aware queries through the Query Initialization module, which encompasses both detection and recognition queries. The Initialization of tracking queries is presented in Sec 3.6. Subsequently, these queries are fed into the task-aware decoder [14] to explicitly capture discriminative and interactive features for text detection, recognition, and tracking simultaneously. Then, a Prompt Queries Generation Module (PQGM) and a Task-aware Adapter to enable the interaction between the hierarchical tasks including word-level and line-level text spotting, as well as video-level text spotting. During this stage of training, most parameters are frozen. Then, multitasking features are learned by the Task-aware Adapter and PQGM. Firstly, we input the prompt into the PQGM for the task to be performed. Then, the PQGM will generate prompt queries and the prompt queries are sent to the Transformer encoder and Task-aware decoder to guide the model to complete the corresponding task. It is worth noting that our method can not only perform image-level cross-domain but also learn video-level cross-domain adaption. Detailed implementations will be provided in the following subsections.
@@ 3.3 Decoder
For the decoder, we adopt the decoder in [14] to enhance the interaction between different tasks. The decoder consists of the vision-language communication module [14] and the Transformer decoder layer [27]. The task-aware queries, including detection queries, recognition queries, and tracking queries. The tracking queries are detailed in Sec. 3.6. The queries are first fed into a vision-language communication module to enable the detection and recognition queries to interact with each other. Then, we adopt intra-group self-attention and inter-group self-attention to further aggregate the features in the Transformer decoder layer. The inter-group self-attention module is utilized to perform the interaction between tracking queries. The inter-group self-attention is used to enable detection and recognition queries to interact with each other.
@@ 3.4 Prompt Queries Generation Module
To enhance the model’s ability to handle multiple tasks effectively, we propose the Prompt Queries Generation Module (PQGM) for generating prompt queries to guide its operation. The structure of this module is illustrated in Fig. 4. For each task, we employ a learnable embedding as the prompt query, with its dimension matching that of the hidden features from the Transformer. Before feeding these prompt queries into the model, we facilitate information exchange between prompt queries for different tasks using an attention mechanism. Subsequently, we input the prompt queries into the model to guide its learning of task-specific features. With the PQGM, VimTS can adeptly handle multiple tasks concurrently and encourage explicit interactions among them, thereby promoting synergies across different tasks. Specifically, following the interaction, we transmit the prompt queries to the Transformer encoder, enabling it to learn task-specific features. These task-specific features are then utilized to assist in Query Initialization and guide the decoder to output the results of corresponding tasks. To further enhance the learning of task-specific features, we further integrate the prompt queries with the task-aware queries from the Query Initialization. We achieve this fusion by adding the prompt queries to the task-aware queries. By the prompt of the prompt queries, the model is guided toward completing the corresponding tasks more effectively.
@@ 3.5 Task-aware Adapter
Taking inspiration from [54], we propose a Task-aware Adapter to dynamically select suitable features for different tasks. The Task-aware Adapter with the PQGM efficiently transforms the original single-task model into a multi-task model suitable for both image and video scenarios, with minimal additional parameters required. The task-aware adapter employs a cascade adapter structure, with one adapter encoding detection information and another encoding recognition information. To implement this, we begin by freezing most parameters of a pre-trained text spotter. Next, we integrate the adapter into the neural network, such as a deformable transformer layer. During multi-task training, the adapter learns the characteristics of different tasks. Notably, the task-aware adapter can not only be applied in image-level scenarios but also learn temporal information, thus aiding the transition of pre-trained models to tasks involving video text spotting. The overall structure is depicted in Fig. 5. Initially, we use two linear layers to reduce the dimension of group queries to $C/4$, thus reducing the parameters of the subsequent module. The task-aware queries then aggregate detection information through an attention mechanism. After aggregating the detection features, for image-level text spotting, an attention mechanism is employed to learn the relationships between different text instances, while for video-level text spotting, the attention mechanism is further used to model the temporal information. The second adapter follows a similar process but focuses on aggregating recognition information. With the task-aware adapter, interactive and discriminative features are efficiently learned, requiring minimal additional parameters. The effectiveness of this module is demonstrated in the ablation study.
@@ 3.6 Tracking Queries
Inspired by MOTR [40] and ColTrack [55], we adopt tracking queries to make the model support tracing text. This approach enables VimTS to dynamically adapt to different inputs including image and video. When presented with the initial frame of a video, the detection queries and recognition queries are used to locate and recognize text instances. These queries are then fed into the decoder, facilitating the aggregation of features and simultaneous decoding of both detection and recognition results. Once the features are aggregated, the semantic and location information of the text instance is encoded into the queries. The same text instance between the front and back frames exhibits a strong correlation, and there is also a strong correlation between their features. Therefore, for subsequent input video frames, we utilize previous detection and recognition queries as input queries, termed tracking queries. These tracking queries are responses to locating, recognizer, and tracking the same text instances in the current frame text instances from the previous frame. In the case of newborn text instances, we continue to employ detection and recognition queries to locate and recognize them. The overall process is shown in Fig 3. The different queries explicitly model the discriminative and interactive features for different tasks in a single decoder and output the detection, recognition, and tracking results simultaneously. The previous end-to-end method [41] employs a relatively separate recognition head after the tacking process. Therefore, the recognition information between the different frames is not fully utilized. By benefiting from a unified framework and incorporating group queries, our method allows for the effective utilization of recognition information across various frames during the tracking process. The recognition information of previous frames can be used to help recognize the text instances in later frames, which can not be achieved by the previous end-to-end method.
@@
MPI-Cooking [60] 24 465K 600s cooking activities middle 0
Hollywood [61] 375 115K 30s human actions in movie middle 0
Breakfast [62] 1.0K 791K 120s breakfast preparation low 0
@@ 3.7 Optimization
Figure 6: Overall framework of CoDeF-based synthetic method.
@@
$\small\mathcal{C}_{match}(Y_{i},\hat{Y}_{\sigma(i)})=\lambda_{c}\mathcal{C}_{c}(\hat{p}_{\sigma(i)}(c_{i}))+\mathds{1}_{\left\{c_{i}\neq\emptyset\right\}}\lambda_{b}\mathcal{C}_{b}(b_{i},\hat{b}_{\sigma(i)})\,,$ (2)
@@
$\small\mathcal{L}(Y_{i},\hat{Y}_{\sigma(i)})=\alpha_{c}\mathcal{L}_{c}(\hat{p}_{\sigma(i)}(c_{i}))+\mathds{1}_{\left\{c_{i}\neq\emptyset\right\}}\mathcal{L}_{b}(b_{i},\hat{b}_{\sigma(i)})\\ +\mathds{1}_{\left\{c_{i}\neq\emptyset\right\}}\alpha_{r}\mathcal{L}_{r}(r_{i},\hat{r}_{\sigma(i)})+\mathds{1}_{\left\{c_{i}\neq\emptyset\right\}}\alpha_{p}\mathcal{L}_{p}(p_{i},\hat{p}_{\sigma(i)})\,,$ (3)
@@ 3.8.1 Data Source and Preprocessing
We manually collect and filter text-free, open-source and unrestricted videos from NExT-QA [56], Charades-Ego [59], Breakfast [62], A2D [57], MPI-Cooking [60], ActorShift [58] and Hollywood [61]. Subsequently, we arrange them in order of priority, which is determined by the stability of transitions, high resolution, and the prevalence of extensive planar areas in the videos. The statistics of each dataset are shown in Tab. I. To achieve distributed processing and GPU memory reduction, we split the videos into segments that contain 368K frames for data synthesis. The synthetic data is termed VTD-368K.
@@ 3.8.2 Synthetic Method
$\small\{\textbf{C},\textbf{D}\}=\mathbb{F}([I_{1},I_{2},...,I_{N}]),$ (4)
@@
$\small p_{k+1}=\mathit{F}_{k\Leftrightarrow k+1}(p_{k})+p_{k},p_{k}\in T_{k},k\in[t,N-1),$ (8)
$\small p_{t-k-1}=\mathit{F}_{t-k\Leftrightarrow t-k-1}(p_{k})+p_{t-k},p_{t-k}\in T_{t-k},k\in[0,t),$ (9)

### v4

@@ Abstract
Unified OCR; Synergy; Video Text Spotting; Text Detection; Text Recognition; Text Tracking
@@ 1 Introduction
Figure 1: Fig. (a) and (b) are two types of cross-domain text spotting, including image-to-image and image-to-video. TT represents the TotalText. IC15 represents the ICDAR2015 [8]. IC13 represent the ICDAR2013 video [9]. TT represents the TotalText [10].
1) In image-level cross-domain text spotting, the challenge lies in the varying styles, fonts, and backgrounds across different benchmarks, requiring models to generalize well beyond their training data. When the training set and the test set belong to different benchmarks, the model usually performs poorly, as demonstrated in  [7]. Furthermore, the differences in annotation formats across various datasets bring further challenges. For instance, TotalText [10] and ICDAR2015 [8] utilize word-level annotations, while CTW1500 [11] employs line-level annotations. Integrating datasets with distinct annotation formats simultaneously proves challenging. This disparity not only complicates the training process but also hampers model performance, as the model struggles to adapt to divergent annotation schemes.
Figure 2: Applying static text spotting methods (TESTR, results shown in the image) to videos, even those with minimal motion, leads to poor performance in both bounding box recall and recognition accuracy.
@@ 2.1 Scene Text Spotting
Scene text spotting has seen a significant shift from conventional methodologies to more modern end-to-end learning techniques. Traditional approaches, as highlighted in previous research [16], often faced challenges related to inference inefficiencies and error accumulation. Pioneering a more unified perspective, Li et al. [17] integrated detection and recognition into a unified framework, creating a comprehensive framework, which predominantly focused on horizontal texts.
@@ 2.2 Video Text Spotting
Compared to static text spotting, there have been few recent advancements in video text spotting due to its challenges. Yin et al. [31] surveyed text detection, tracking, and recognition in video. Wang et al. [32] developed an end-to-end system for video text spotting, utilizing hand-crafted post-processing for tracking. Subsequently, Cheng et al. [33] integrated a spatial-temporal video text detector with a text recommender to select high-quality text and perform recognition once. Its improved version, FREE [34], adopted an end-to-end trainable framework for faster and robust video text spotting. Nguyen et al. [35] performed multi-scale character detection and exploited temporal redundancy to remove false positives. Rong et al. [36] proposed a framework for scene text recognition across multiple frames using feature representation and a conditional random field model. Wu et al. [37] developed a dual decoder framework for detection and tracking, utilizing object features and IoU match. Inspired by the MOTR [38], Wu et al. [39] unified text detection and tracking into a single decoder, treating video text spotting as a long-range temporal variation problem. To further enhance the performance of the recognizer, VLSpotter [40] introduced a super-resolution module and language model for video text spotting.
@@ 2.3 Domain Adaptation
Domain adaptation, a burgeoning field of study, addresses the challenges posed by domain shift and dataset bias. It has been studied extensively in many fields [41, 42, 43]. Researchers also proposed several cross-domain approaches in OCR. For instance, Chen et al. [44] propose a pixel-level domain adaptation block and an image-level domain adaptation block for text detection to reduce the domain shift by introducing adversarial training. Wu et al. [45] attempt to transfer knowledge from synthetic data to real data for scene text detection. They employ text instance alignment to facilitate learning domain-invariant features adversarially, and utilize text self-training to reduce the impact of inaccurate pseudo-labels. Additionally, Chen et al. [46] develop a self-training framework aimed at automatically extracting the pseudo-labels of challenging text instances from unannotated videos or images. Yu et al. [47] introduce a cross-modal interaction mechanism, integrating vision-language information from CLIP [48] to improve the cross-domain performance of scene text detectors. The above methods are designed for scene text detection.
For scene text recognition, Zhang et al. [49] introduced SSDAN, a sequence-to-sequence domain adaptation network. SSDAN utilizes a gated attention similarity unit to dynamically align the distribution of source and target sequence data within a character-level feature space. GA-DAN [50] develops a geometry-aware domain adaptation network for both text detection and recognition. It effectively handles cross-domain shifts by simultaneously modeling changes in both geometric and appearance spaces, enabling realistic image conversion between domains with diverse characteristics. E2STR [51] presents an in-context training strategy for scene text recognition, which enables text recognizer to perform rapid adaptation across diverse scenarios without additional fine-tuning. In recent years, there have also been ways to start looking at cross-domain aspects of text spotting. Yu et al. [7] build on their prior work [47] by introducing FastTCM-CR50, a new backbone for scene text detection and spotting that improves cross-domain performance. It integrates an implicit image condition and bimodal similarity matching, enabling offline CLIP computations to boost accuracy and reduce inference time.
@@ 3 Methodology
In this paper, we introduce VimTS, a unified framework designed to leverage the synergy between various tasks. Through this unified framework, this approach aims to improve the generalization ability of the text spotter. The overall architecture is illustrated in Fig. 3. Inspired by [14], we employ a set of task-aware queries to represent various tasks. Initially, image features are obtained via a feature extraction process comprising ResNet50, the receptive enhancement module (REM), and the Transformer encoder. Similar to [14], these features are then used to generate task-aware queries through the Query Initialization module, which encompasses both detection and recognition queries. The Initialization of tracking queries is presented in Sec. 3.6. Subsequently, these queries are fed into the task-aware decoder [14] to explicitly capture discriminative and interactive features for text detection, recognition, and tracking simultaneously. Then, a Prompt Queries Generation Module (PQGM) and a Task-aware Adapter are used to enable the interaction between the hierarchical tasks including word-level and line-level text spotting, as well as video-level text spotting. During this stage of training, most parameters are frozen. Then, multitasking features are learned by the Task-aware Adapter and PQGM. Firstly, we input the prompt into the PQGM for the task to be performed. Then, the PQGM will generate prompt queries and the prompt queries are sent to the Transformer encoder and Task-aware decoder to guide the model to complete the corresponding task. It is worth noting that our method can not only perform image-level cross-domain but also learn video-level cross-domain adaption. Detailed implementations will be provided in the following subsections.
@@ 3.3 Decoder
For the decoder, we adopt the decoder in [14] to enhance the interaction between different tasks. The decoder consists of the vision-language communication module and the Transformer decoder layer [27]. The task-aware queries, including detection queries, recognition queries, and tracking queries. The tracking queries are detailed in Sec. 3.6. The queries are first fed into a vision-language communication module to enable the detection and recognition queries to interact with each other. Then, we adopt intra-group self-attention and inter-group self-attention to further aggregate the features in the Transformer decoder layer. The inter-group self-attention module is utilized to perform the interaction between tracking queries. The inter-group self-attention is used to enable detection and recognition queries to interact with each other.
@@ 3.4 Prompt Queries Generation Module
To enhance the model’s ability to handle multiple tasks effectively, we propose the Prompt Queries Generation Module (PQGM) for generating prompt queries to guide its operation. The structure of this module is illustrated in Fig. 4. When inputting a prompt for a task, such as word-level text spotting, we use an embedding layer to transform the task prompt into prompt queries. These prompt queries then interact with queries from other tasks via an attention mechanism. Following the interaction, we integrate the prompt queries with the image features by directly adding them together. The combined features are then fed into the Transformer encoder for further processing and fusion. To adapt these prompt queries to learn the task-specific features, we adopt the adapter [53] in the Transformer encoder. This adapter consists of two linear projection layers and a nonlinearity activation function. These task-specific features are then utilized in Query Initialization and the Transformer decoder. To further enhance the learning of task-specific features, we further add the prompt queries to the task-aware queries from the Query Initialization. By the prompt of the prompt queries, the model is guided toward completing the corresponding tasks more effectively.
@@ 3.5 Task-aware Adapter
Inspired by [53], we propose a Task-aware Adapter that dynamically selects features tailored for different tasks. Combined with the PQGM, this adapter transforms the original single-task model into a multi-task model capable of handling both image and video scenarios, with minimal additional parameters. The adapter uses a cascade structure: one adapter encodes detection information, while another encodes recognition. To implement this, we freeze most parameters of a pre-trained text spotter and integrate the adapter into the neural network, such as a deformable transformer layer. During multi-task training, the adapter learns to adapt to the characteristics of each task. The task-aware adapter is effective not only in image-level tasks but also in learning temporal information, enabling the transition of pre-trained models to video text spotting tasks. As shown in Fig. 5, we first use two linear layers to reduce the dimensionality of group queries to $C/4$, minimizing the parameters of the subsequent module. Task-aware queries then aggregate detection features via an attention mechanism. For image-level spotting, the attention learns relationships between text instances, while for video-level tasks, it captures temporal information. The second adapter follows a similar process, focusing on recognition. This approach efficiently learns interactive and discriminative features with minimal additional parameters, as validated in the ablation study (Sec. 4.3).
@@ 3.6 Tracking Queries
Inspired by MOTR [38] and ColTrack [54], we incorporate tracking queries to enable the model to trace text. This approach enables VimTS to adapt dynamically to both images and videos. For the initial video frame, detection and recognition queries are used to locate and recognize text instances, which are then fed into the decoder to aggregate features and simultaneously decode detection and recognition results. Once aggregated, the queries encode both semantic and location information. Since text instances across consecutive frames are strongly correlated, tracking queries, based on prior frame detection and recognition results, are used for subsequent frames to track and recognize the same instances. For new text instances, detection and recognition queries are used as usual.
The overall process is shown in Fig 3. The different queries explicitly model the discriminative and interactive features for different tasks in a single decoder and output the detection, recognition, and tracking results simultaneously. The previous end-to-end method [39] employs a relatively separate recognition head after the tacking process. Therefore, the recognition information between the different frames is not fully utilized. By benefiting from a unified framework and incorporating group queries, our method allows for the effective utilization of recognition information across various frames during the tracking process. The recognition information of previous frames can be used to help recognize the text instances in later frames, which can not be achieved by the previous end-to-end method.
@@ 3.7 Optimization
Figure 6: Overall framework of the proposed synthetic method.
@@
$\mathcal{C}_{match}(Y_{i},\hat{Y}_{\sigma(i)})=\lambda_{c}\mathcal{C}_{c}(\hat{p}_{\sigma(i)}(c_{i}))+\mathds{1}_{\left\{c_{i}\neq\emptyset\right\}}\lambda_{b}\mathcal{C}_{b}(b_{i},\hat{b}_{\sigma(i)})\,,$ (2)
@@
$\mathcal{L}(Y_{i},\hat{Y}_{\sigma(i)})=\alpha_{c}\mathcal{L}_{c}(\hat{p}_{\sigma(i)}(c_{i}))+\mathds{1}_{\left\{c_{i}\neq\emptyset\right\}}\mathcal{L}_{b}(b_{i},\hat{b}_{\sigma(i)})\\ +\mathds{1}_{\left\{c_{i}\neq\emptyset\right\}}\alpha_{r}\mathcal{L}_{r}(r_{i},\hat{r}_{\sigma(i)})+\mathds{1}_{\left\{c_{i}\neq\emptyset\right\}}\alpha_{p}\mathcal{L}_{p}(p_{i},\hat{p}_{\sigma(i)})\,,$ (3)
@@ 3.8.1 Data Source and Preprocessing
The videos used for synthesizing data must not contain text of their own and publicly available without copyright issues. Additionally, the video should also be relatively clear, ensuring the quality of the synthesized dataset. According to this requirement, we carefully selected the video dataset used in the paper from publicly available datasets.
We exclude text-containing video datasets and manually filter single-thesis, clustered-scene, and low-resolution public video datasets. We found that the unsuitable video dataset for synthesis can be broadly categorized into four categories: 1) Text-Containing Datasets with labeling the text: These include OTB2015 [64], ActivityNet-Caption [65], and AVA [66]. The presence of unlabeled text in these videos can confuse the model during the training phase when it encounters synthesized data. 2) Single-Theme Datasets: Examples such as EPIC-KITCHENS [67] and MPI-Cooking [68] lack the thematic diversity required to improve the model’s generalization capabilities. 3) Cluttered-Scene Datasets: Videos from collections like Hollywood [69] and Breakfast [70] do not offer clear areas suitable for rendering, due to their complex and cluttered scenes. 4) Low-Resolution Datasets: Such as TinyVIRAT [71], these datasets fail to provide the necessary visual clarity and detail for high-quality synthesis. Following this evaluation, we selected a subset of text-free, open-source, and unrestricted videos from NExT-QA [55], A2D [56], ActorShift [57], and Charades-Ego [58].
The statistics of each dataset are shown in Tab. I. To achieve distributed processing and GPU memory reduction, we split the videos into segments that contain 368k frames for data synthesis. The synthetic data is termed VTD-368k.
In the VTD-368k dataset, the average text length is 4.14 characters, with an average of 53.45 characters appearing per frame. We define text density as the ratio of the number of text instances to the total number of frames. According to this definition, the text density of VTD-368k is 6.44. Additionally, the text styles within VTD-368k are kept consistent with those found in SynthText [13].
@@ 3.8.2 Synthetic Method
$\{\textbf{C},\textbf{D}\}=\mathbb{F}([I_{1},I_{2},...,I_{N}]),$ (4)
@@
$p_{k+1}=\mathit{F}_{k\Leftrightarrow k+1}(p_{k})+p_{k},p_{k}\in T_{k},k\in[t,N-1),$ (8)
$p_{t-k-1}=\mathit{F}_{t-k\Leftrightarrow t-k-1}(p_{k})+p_{t-k},p_{t-k}\in T_{t-k},k\in[0,t),$ (9)